# FedBiP: Heterogeneous One-Shot Federated Learning with Personalized Latent Diffusion Models

### v1

@@ 5.2 Validation on Medical and Satellite Image Datasets
To illustrate the effectiveness of FedBiP on challenging real-world applications, we adopt a medical dataset, DermaMNIST (Yang et al., 2023a), comprising dermatoscopic images of 7 types of skin lesion, and a satellite image dataset, UC Merced Land Use Dataset (UCM) (Yang & Newsam, 2010), which includes satellite images representing 21 different land use categories. We assume there are 5 research institutions (clients) participating in the collaborative training. To construct local datasets for each client in OSFL, we employ the Dirichlet distribution $Dir_{\beta}$ to model label space heterogeneity, in which a smaller $\beta$ indicates higher data heterogeneity. Following (Zhou et al., 2022), we use the textual template ”A dermatoscopic image of a [CLS], a type of pigmented skin lesions.” and ”A centered satellite photo of [CLS].” for DermaMNIST and UCM, respectively.
@@ 5.3 Ablation Study
To illustrate the importance of different FedBiP components, we conduct an ablation study on three OSFL benchmark datasets. The results are shown in Table 5.3. First, we observe that simply prompting LDM with ”A [STY] style of a [CLS]” and synthesizing images at central server is ineffective. Next, we notice that optimizing only the category concept vector $V_{c}$ leads to only minimal performance improvements. We hypothesize that this is because the categories in these benchmarks are general objects, such as ”person” or ”clock”, which are already well-captured by LDM during pretraining. In contrast, optimiz-
Table 3: Ablation study for different components of FedBiP on three benchmarks.
@@
ing the domain concept vector $S$ produces visible performance gain. This can be attributed to the mismatch between the textual representation of domain concepts and LDM’s pretraining. For example, as described in Motivation section (Figure 1), ”Quickdraw” in DomainNet encompasses images characterized by very simple lines, while LDM tends to generate images with finer details. Furthermore, applying instance-level personalization with $z(T)$ yields a performance boost, highlighting the importance of fine-grained personalization in improving LDM. Finally, combining both levels of personalization further improves the results, which demonstrates their complementarity.
@@ 5.4 Scalability Analysis of FedBiP
The validation results are presented in Figure 5.5. We observe that the performance of the baseline method FedAvg remains unchanged with the addition of more clients to FL. In contrast, the validation performance of FedBiP consistently increases, narrowing the gap between distributed opti-
Figure 4: Validation results of FedBiP with varying number of clients on DomainNet.
Figure 5: Validation results of FedBiP with synthesizing different numbers of images at central server.
mization and Central optimization. Furthermore, FedBiP outperforms FedAvg by $9.51\%$ when the largest number of clients join FL, further indicating its scalability for real-world complex federated systems with more clients.
@@ 5.5 Privacy Analysis
Pixel Value Histogram Analysis: To further analyze FedBiP from a statistical perspective, we provide histograms of both $\overline{z}(0)$ (the interpolated latent vectors of input images) and the corresponding $z(T)$ ($\overline{z}(0)$ with $T$-steps of random noise added). We observe that $z(T)$ closely resembles a standard Gaussian distribution, which contains less information about the original input images compared to $\overline{z}(0)$. This indicates that transmitting the noised $z(T)$ is more private than $\overline{z}(0)$, and would not significantly compromise privacy regulations. Additionally, we notice that $\overline{z}(0)$ could be further replaced with the average latent vectors of all samples from a specific class, i.e., categorical prototypes (Tan et al., 2022). This substitution might further protect client privacy and is appropriate for applications with stringent privacy requirements. We leave this for future work.
Membership Inference Attack (MIA) Analysis: Finally, we analyze the resilience of FedBiP against MIA. Following (Yeom et al., 2018; Salem et al., 2018), we compute the average loss and entropy of the final model on both training member and non-member data, and report the difference between the two averages. A smaller difference corresponds to better membership privacy preservation. From the MIA Analysis in Table 5.5, we can observe that FedBiP demonstrates superior resilience against MIA.
@@ 5.7 Visualization for Challenging Domains
In this section, we present the synthetic images generated for the challenging domains, i.e., Quickdraw (DomainNet) and Sketch (PACS), as shown in Figure 8. Our observations indicate that FedBiP achieves superior generation quality by more accurately adhering to the original distribution of clients’ local data compared to the diffusion-based method FGL (Zhang et al., 2023). This visualization further highlights the effectiveness of our bi-level personalization approach.
Figure 8: Comparison of synthetic images for challenging domains.
@@ 6 Conclusion
In this work, we propose the first framework to address feature space heterogeneity in One-Shot Federated Learning (OSFL) using generative foundation models, specifically Latent Diffusion Model (LDM). The proposed method, named FedBiP, personalizes the pretrained LDM at both instance-level and concept-level. This design enables LDM to synthesize images that adhere to the local data distribution of each client, exhibiting significant deviations compared to its pretraining dataset. The experimental results indicate its effectiveness under OSFL systems with both feature and label space heterogeneity, surpassing the baseline and multiple concurrent methods. Additional experiments with medical or satellite images demonstrate its maturity for challenging real-world applications. Moreover, additional analysis highlights its promising scalability and privacy-preserving capability.
Appendix A Experimental Details
We use 1 NVIDIA RTX A5000 with 24GB RAM to run the experiments. We use PyTorch (Paszke et al., 2019) to implement our algorithm. For the baseline FedAvg, the total communication round is set to 50. For FGL (Zhang et al., 2023), we generate 3500 images per class per domain. For the optimization of the classification model, we use SGD with momentum as the optimizer, where the learning rate is set to 0.01 and the momentum is 0.9. The optimization epoch is set to 50. The training image resolution is set to $512\times 512$ for all datasets.
For FedD3 (Song et al., 2023), we adopt Kernel Inducing Points (KIP) to distill the original dataset into 1 image per class per domain and transmit them to the central server. For DENSE (Zhang et al., 2022), we first finetune the pretrained ResNet-18 (He et al., 2016) at each client and then optimize a Generator to conduct model distillation at central server. The hyperparameters used in these methods are following their original papers. For FedMLA, we use Adam optimizer to optimize the concept vectors. The learning rate is set to 0.1 and beta is set to (0.9, 0.999). The total training epochs is set to 30. We adopt the Pseudo Numerical Diffusion Model (PNDM) (Liu et al., 2022) in the Latent Diffusion Model. The perturbation intensity for domain concept vector $\sigma_{\mu}$ is set to 0.1 for all dataset. More dataset specific hyperparameters are provided in Table 5.
Table 5: Detailed hyperparameters for each dataset. The highlighted words ([STY]) in the textual prompt will be replaced by the domain concept vectors. The [CLS] will be replaced by the class concept vectors.
Dataset prompt $n_{s}$ $n_{c}$ $C$ Class Names
Derma MNIST A dermatoscopic image of a [CLS], a type of pigmented skin lesions. 2 4 10 intraepithelial carcinoma, basal cell carcinoma, benign keratosis, dermatofibroma, melanoma, melanocytic nevi, vascular skin
UCM A centered satellite photo of [CLS]. 3 3 21 agricultural, dense residential, medium residential, sparse residential, parking lot, buildings, harbor, mobile homepark, storage tanks, freeway, intersection, overpass, golf course, baseball diamond, runway, tenniscourt, beach, forest, river, chaparral, airplane
Domain Net A [STY] of [CLS]. 1 1 10 airplane, clock, axe, basketball, bicycle, bird, strawberry, flower, pizza, bracelet
Office Home A [STY] of [CLS]. 1 1 20 Marker, Spoon, Pencil, Speaker, Toys, Fan, Hammer, Notebook, Telephone, Sink, Chair, Fork, Kettle, Bucket, Knives, Monitor, Mop, Oven, Pen, Couch
PACS A [STY] of [CLS]. 1 1 7 dog, elephant, giraffe, guitar, horse, house, person
Appendix B Synthetic Image Visualization
We provide synthetic images for all benchmarks in the following figures, where we observe that the synthetic images generally follow the distribution and characteristics of the original training datasets at each client. Besides, the visual quality of the generated images, e.g., the detailed features of the objects, is also promising.
Figure 9: Synthetic Images for DomainNet benchmark.
Figure 10: Synthetic Images for PACS benchmark.
Figure 11: Synthetic Images for OfficeHome benchmark.
Figure 12: Synthetic Images for UCM benchmark.
Figure 13: Synthetic Images for DermaMNIST benchmark.

### v2

@@ 5.1 Benchmark Experiments
5.2 Validation on Medical and Satellite Image Datasets
To illustrate the effectiveness of FedBiP on challenging real-world applications, we adopt a medical dataset, DermaMNIST [42], comprising dermatoscopic images of 7 types of skin lesion, and a satellite image dataset, UC Merced Land Use Dataset (UCM) [46], which includes satellite images representing 21 different land use categories. We assume there are 5 research institutions (clients) participating in the collaborative training. To construct local datasets for each client in OSFL, we employ the Dirichlet distribution $Dir_{\beta}$ to model label space heterogeneity, in which a smaller $\beta$ indicates higher data heterogeneity. Following [52], we use the textual template ”A dermatoscopic image of a [CLS], a type of pigmented skin lesions.” and ”A centered satellite photo of [CLS].” for DermaMNIST and UCM, respectively.
@@ 5.3 Ablation Study
To illustrate the importance of different FedBiP components, we conduct an ablation study on three OSFL benchmark datasets. The results are shown in Table 3. First, we observe that simply prompting LDM with ”A [STY] style of a [CLS]” and synthesizing images at central server is ineffective. Next, we notice that optimizing only the category concept vector $V_{c}$ leads to only minimal performance improvements. We hypothesize that this is because the categories in these benchmarks are general objects, such as ”person” or ”clock”, which are already well-captured by LDM during pretraining. In contrast, optimizing the domain concept vector $S$ produces visible performance gain. This can be attributed to the mismatch between the textual representation of domain concepts and LDM’s pretraining. For example, as described in Motivation section (Figure 1), ”Quickdraw” in DomainNet encompasses images characterized by very simple lines, while LDM tends to generate images with finer details. Furthermore, applying instance-level personalization with $z(T)$ yields a performance boost, highlighting the importance of fine-grained personalization in improving LDM. Finally, combining both levels of personalization further improves the results, which demonstrates their complementarity.
Table 3: Ablation study for different components of FedBiP.
@@ 5.4 Scalability Analysis of FedBiP
The validation results are presented in Figure 3. We observe that the performance of the baseline method FedAvg remains unchanged with the addition of more clients to FL. In contrast, the validation performance of FedBiP consistently increases, narrowing the gap between distributed optimization and Central optimization. Furthermore, FedBiP outperforms FedAvg by $9.51\%$ when the largest number of clients join FL, further indicating its scalability for real-world complex federated systems with more clients.
@@
Figure 3: Validation results with varying client numbers on DomainNet.
Figure 4: Validation results with synthesizing different numbers of images at central server.
@@ 5.5 Privacy Analysis
Pixel Value Histogram Analysis: To further analyze FedBiP from a statistical perspective, we provide histograms of both $\overline{z}(0)$ (the interpolated latent vectors of input images) and the corresponding $z(T)$ ($\overline{z}(0)$ with $T$-steps of random noise added). We observe that $z(T)$ closely resembles a standard Gaussian distribution, which contains less information about the original input images compared to $\overline{z}(0)$. This indicates that transmitting the noised $z(T)$ is more private than $\overline{z}(0)$, and would not significantly compromise privacy regulations. Additionally, we notice that $\overline{z}(0)$ could be further replaced with the average latent vectors of all samples from a specific class, i.e., categorical prototypes [40]. This substitution might further protect client privacy and is appropriate for applications with stringent privacy requirements. We leave this for future work.
Membership Inference Attack (MIA) Analysis: Finally, we analyze the resilience of FedBiP against MIA. Following [47, 34], we compute the average loss and entropy of the final model on both training member and non-member data, and report the difference between the two averages. A smaller difference corresponds to better membership privacy preservation. From the MIA Analysis in Table 4, we can observe that FedBiP demonstrates superior resilience against MIA.
@@ 5.7 Visualization for Challenging Domains
In this section, we present the synthetic images generated for the challenging domains, i.e., Quickdraw (DomainNet) and Sketch (PACS), as shown in Figure 7. Our observations indicate that FedBiP achieves superior generation quality by more accurately adhering to the original distribution of clients’ local data compared to the diffusion-based method FGL [51]. This visualization further highlights the effectiveness of our bi-level personalization approach.
Figure 7: Comparison of synthetic images from different algorithms for challenging domains.
@@ 6 Conclusion
In this work, we propose the first framework to address feature space heterogeneity in One-Shot Federated Learning (OSFL) using generative foundation models, specifically Latent Diffusion Model (LDM). The proposed method, named FedBiP, personalizes the pretrained LDM at both instance-level and concept-level. This design enables LDM to synthesize images that adhere to the local data distribution of each client, exhibiting significant deviations compared to its pretraining dataset. The experimental results indicate its effectiveness under OSFL systems with both feature and label space heterogeneity, surpassing the baseline and multiple concurrent methods. Additional experiments with medical or satellite images demonstrate its maturity for challenging real-world applications. Moreover, additional analysis highlights its promising scalability for more complex OSFL systems and privacy-preserving capability against different types of attacks.